answer all parts.

A stock has a beta of 1.26 and an expected return of 12.4 percent. A risk-free asset currently earns 4.1 percent.

a. What is the expected return on a portfolio that is equally invested in the two assets? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Expected return ____ % ?

b. If a portfolio of the two assets has a beta of .86, what are the portfolio weights? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)

Weight of the stock
Weight of the risk-free asset

??

c. If a portfolio of the two assets has an expected return of 11.6 percent, what is its beta? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Beta ____?   

d. If a portfolio of the two assets has a beta of 2.46, what are the portfolio weights? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)

Weight of the stock
Weight of the risk-free asset

???
Answer #1

a.

Expected return

=50%*12.4%+50%*4.1%

=8.25%

b.

Weight of the stock=0.86/1.26=0.6825 or 68.2540%

Weight of the risk-free asset=1-0.6825=0.3175 or 31.7460%

c.

assume weight of stock=S

risk free rate weight=1-S

Expected return=11.6%

S*12.4%+(1-S)*4.1%=11.6%

S*(12.4%-4.1%)=11.6%-4.1%

S=(11.6%-4.1%)/(12.4%-4.1%)

=0.9036 or 90.3614%

what is its beta=90.3614%*1.26=1.14

d.

weight of stock*beta of stock+weight of risk free rate*beta of risk free rate=2.46

S*1.26+(1-S)*0=2.46

S=2.46/1.26

=195.2381% OR 1.9524 is Weight of the stock

Weight of the risk-free asset=1-195.2381%=-95.2381% or -0.9524

the above is answer..